If the rate of production of HCl is 3.0 X 10-5 M/s at a particular instant. What is the rate of consumption of SnCl4 at this same time?
1
Expert's answer
2020-04-20T15:08:56-0400

Let's write the balanced reaction equation:


SnCl4 + 4H2O "\rightarrow" Sn(OH)4 + 4HCl


As you can see from this equation, when 1 mol of SnCl4 reacts, 4 mol of HCl are produced. The rates of the reagents consumption and the rates of the products formation are related by stoichiometric coefficients. The rate of the reaction at a particular instant can be expressed as:

"r = -\frac{\text{d}[\text{SnCl}_4]}{\text{d}t} =\frac{1}{4}\cdot \frac{\text{d}[\text{HCl}]}{\text{d}t}",

where "\frac{\text{d}[\text{SnCl}_4]}{\text{d}t}"  is the rate of consumption of SnCl4 and "\frac{\text{d}[\text{HCl}]}{\text{d}t}" is the rate of formation of HCl. The of consumption of SnCl4 at the same time is:

"\frac{\text{d}[\text{SnCl}_4]}{\text{d}t} =- \frac{1}{4}\cdot 3.0 \cdot 10^{-5} =- 0.75\cdot 10^{-5}" M/s

Answer: if the rate of production of HCl is 3.0 · 10-5 M/s at a particular instant, the rate of consumption of SnCl4 at this same time is -0.75 ·10-5 M/s